# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v1

@@ Abstract
In this paper, we propose a method for transferring feature representation to lightweight student models from larger teacher models. We mathematically define a new notion called perception coherence. Based on this notion, we propose a loss function, which takes into account the dissimilarities between data points in feature space through their ranking. At a high level, by minimizing this loss function, the student model learns to mimic how the teacher model perceives inputs. More precisely, our method is motivated by the fact that the representational capacity of the student model is weaker than the teacher model. Hence, we aim to develop a new method allowing for a better relaxation. This means that, the student model does not need to preserve the absolute geometry of the teacher one, while preserving global coherence through dissimilarity ranking. Our theoretical insights provide a probabilistic perspective on the process of feature representation transfer. Our experiments results show that our method outperforms or achieves on-par performance compared to strong baseline methods for representation transferring.
@@ 1 Introduction
The need of lightweight models and Knowledge distilling (KD). Deep learning models are at the forefront of performance in different domains and tasks, such as classification he2016deep ; tan2019efficientnet ; tan2021efficientnetv2 ; he2020momentum or object detection ren2015faster ; lin2017feature ; lin2017focal ; he2020momentum . However, as the performance of neural networks increases, so does the cost of increasingly large models. In many applications where resources are limited (e.g. mobile devices) or one needs a fast execution, it is preferable to use lightweight models. Among the many approaches to obtain lightweight models, KD buciluǎ2006model ; hinton2015distilling ; wang2021knowledge stands out as a particularly interesting direction. A KD setup essentially consists of a teacher model and a student model. The student model (here the targeted lightweight model) is generally smaller than the teacher model, and therefore more efficient for the task in question. At a high level, the KD process consists in teaching the student model so that it mimics the way the teacher model understands the input. This is usually achieved by using a so-called transfer dataset. The student model is trained to behave like the teacher model on this set in some way . Many previous works have shown that KD improves the generalization capacity of the student model hinton2015distilling ; chen2017learning ; zhu2018knowledge ; beyer2022knowledge ; huang2022knowledge . Different up-to-date KD approaches will be discussed in Section 2.
@@
Feature representation transferring. Classical KD techniques generally focus on matching the outputs of student and teacher models, typically for classification tasks hinton2015distilling . However, this approach has limitations, such as requiring the same number of classes between the teacher and student models. To address these issues, previous works passalis2018learning ; passalis2020probabilistic ; tiancontrastive propose to directly transfer the feature space representations through learning. This enables the student model to better capture the input-input relations represented in the feature space of the teacher model, by preserving its geometry and distribution. This leads to more effective knowledge transfer. Furthermore, note that this approach is an unsupervised transferring method, as it only uses a transfer set without labels. As a consequence, the learned features can be used for various downstream tasks. Because of these appealing properties, our method aligns with this last approach.
@@
Motivation and an overview of our method. Generally, the student model is much smaller than the teacher model. Thus, in terms of representational power, the student model cannot learn to produce features with the exact same geometry as the teacher model. Hence, we aim to develop a method such that (1) the student model does not need to copy the feature representation of the teacher model; (2) the learned representation retains a certain overall consistency with the teacher model’s feature representation and (3) the method works for representation transferring between spaces having different dimensions. For these purposes, we develop a novel fairly simple notion called perception coherence, that is explained subsequently. To begin with, let us consider a reference input $x$ and a set of $n$ point $\{x_{i}\}_{i}^{n}$ (called compared set). These points are fed into the student and teacher models and are mapped into the feature space of each model. The feature space reflects how the models perceive the inputs. Placing ourselves in the two feature spaces, let $d_{1}$ and $d_{2}$ be dissimilarity measures for the teacher and student models, respectively. Then, if $d_{1}(x,x_{i})\leq d_{1}(x,x_{j})$ for some $i,j\in\{1,2,\dotsc,n\}$ ($i\neq j$), we expect that $d_{2}(x,x_{i})\leq d_{2}(x,x_{j})$ also holds. Less formally, if the teacher model perceives $x_{i}$ to be more similar to $x$ than $x_{j}$, we expect the student to have the same perception. We refer to this as the property of perception coherence. By adopting this behavior for different reference points and compared sets, the student model learns to have a coherent perception with respect to the teacher model. The general idea of our method is illustrated in Figure 1.1. Note that we do not measure how much more similar $x_{i}$ is to $x$ than $x_{j}$, but we only rank the dissimilarities. This allows flexibility, as the student does not need to match the teacher’s geometry or distribution exactly. However, the ranking operation is non-differentiable. Hence, we introduce a simple differentiable soft ranking function (Eq. (3.3)). Our contribution can be summarized as follows:
  • We propose a new method for distilling knowledge to lightweight models, based on deep representation transferring. To do so, we develop a novel probabilistic concept called perception coherence. This leads to a novel easy-to-implement loss function (Eq. (3.4)), based on a soft ranking operation.
  • Our theoretical insights via a probabilistic framework provide a new way of understanding representation transfer between models through local and global dissimilarity ranking in the feature spaces between transfer domains. More particularly, we show that a larger global expected coherence level leads to a higher chance that the learned representation respects the dissimilarity ranking of the teacher one.
  • We conduct experiments on retrieval and classification tasks to compare our method with strong baseline methods. The results show that our method outperforms or achieves competitive performance with state-of-the-art methods on lightweight representation transferring settings. Our code is available at anonymized-code.
@@
Figure 1.1: General scheme of our method. In each mini-batch, we use each input as the reference point to compute the dissimilarity to the others (compared set). Then, using loss function in Eq. (3.4), the student model is trained to respect the perception coherence. Notice that $f_{1}$ and $f_{2}$ can be the whole model (when using the output layer) or a part of the model (when using an intermediate layer).
@@ 2 Related Work
KD based on soft label produced by teacher model. This approach has been studied since the early days of neural networks buciluǎ2006model . Therein, the authors show that knowledge can be transferred from a large ensemble of models to a single small model. A standard KD method consists in using output of classification models hinton2015distilling ; tang2016recurrent ; huang2022knowledge , with a temperature, leading to smoother soft labels. Based on the same spirit, chan2015transferring proposes to match the output distributions using Kullback-Leibler divergence. Soft label was also used in the context of domain adaption tzeng2015simultaneous . While all these methods show the benefits of soft labels, the main limit of this approach is the need of having the same number of classes between models.
@@
KD based on distance matching. In romero2014fitnets , the authors propose to minimize the distance between student and teacher’s features to guide the student model during training. However, in the case of dimension mismatch between the models, one need to pass through a linear transformation, which leads to substantial information loss. In yim2017gift , the authors propose to minimize distance of the so-call FSP matrix between teacher and student models (capturing inter-layer relations). More advanced methods consist in minimizing the distance between the derivatives w.r.t. to the features czarnecki2017sobolev or between the attention map of student and teacher zagoruyko2016paying . All these approaches require the dimension match, and need some form of interpolation in the case of dimension mismatch. Overcoming the drawback of dimension mismatch, passalis2018unsupervised proposes to match the the pair-wise similarity matrix of the training points (in the transfer set) between the student and teacher models. The method proposed in yu2019learning is to match the distance between pairs of samples. However, this brute force approach is not likely to work as the student model is much smaller than the teacher one (as discussed in passalis2020probabilistic ).
@@
Feature representation transfer. Some previous works treat the representation transfer as the problem of distribution matching, based on Maximum Mean Discrepancy huang2017like or adversarial training belagiannis2018adversarial . The latter method is inspired by Generative Adversarial Networks (GANs) goodfellow2014generative . One major drawback of distribution matching is the need of having the same dimension between the two feature spaces. Probabilistic frameworks based on mutual information are proposed in ahn2019variational ; tiancontrastive . However, the method in ahn2019variational requires a Gaussian assumption (this is not necessarily satisfied in practice) while the method tiancontrastive requires the input features to have the same dimension (hence requiring a linear transformation, which may reduce the useful information). Overcoming the requirement of dimension match, park2019relational proposes to transfer the input relations using distance-wise and angle-wise distillation loss. However, the structural information is quite limited, as it only considers a single triple of points at a time. Overcoming this drawback, passalis2018learning proposes a method of probabilistic knowledge transfer (PKT). The high level idea of this last method is to transfer the local geometry learned by the teacher model based on kernel techniques. Based on this work, passalis2020heterogeneous ; passalis2020probabilistic basically share the same idea, but they apply the technique for multiples intermediate layers with an auxiliary model. The benefits of using auxiliary models are discussed in mirzadeh2020improved . As in passalis2018learning ; passalis2020heterogeneous ; passalis2020probabilistic , our method takes into account the relation between the inputs and allows representation transfer between space having different dimensions. However, in passalis2018learning ; passalis2020heterogeneous ; passalis2020probabilistic , a common kernel is used for all points (e.g. a Gaussian one with a fixed bandwidth), which cannot adapt to local density variation. This contrasts with our method, where we only rank the dissimilarities but do not use the dissimilarity magnitude. Thus, our method overcomes the problem of local scale variations, allowing itself to adapt flexibly to the local density.
@@ 3.1 Perception coherence and other notions
At a high level, if $f_{1}$ perceives $x_{1}$ to be more similar to $x$ than $x_{2}$, then it should be also the case for $f_{2}$ after the knowledge distilling. This is represented through the feature space of each model, endowed with their corresponding dissimilarity metric. We present this intuition via a more formal concept named absolute perception coherence as follows.
Definition 1 (Absolute perception coherence).
@@
$F_{1}(x,x^{\prime}):=\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(x,x^{\prime})\right),\;F_{2}(x,x^{\prime}):=\mathbb{P}_{X}(d_{2}(x,X)\leq d_{2}(x,x^{\prime}))\ .$ (3.1)
Remark 1.
@@
Remark 2 (Relation between cumulative function and ranking).
Let us consider a discrete distribution composed of $N+1$ examples $\{x\}\bigcup\{x_{i}\}_{i=1}^{N}$ and the extractor $f$ endowed with a dissimilarity metric $d$. We consider $x$ as the reference point and we compute dissimilarity between $x$ and other points $x_{i}$’s to obtain $D:=\{d(x,x_{i})\}_{i=1}^{N}$. Assuming there is no ties, i.e., $d(x,x_{i})\neq d(x,x_{j})$ for all $i\neq j$, then, the cumulative function $F(x,x_{i})$ is the normalized rank of $d(x,x_{i})$ in the set $D$.
Using the cumulative function of Eq. (3.1), we define the perception coherence level as follows.
Definition 2 (Perception coherence level).
@@
$\phi_{f_{1},f_{2}}(x):=1-\mathbb{E}_{X}\left[\left|F_{1}(x,X)-F_{2}(x,X)\right|\right].$ (3.2)
Remark 3.
@@
The proof in this proposition can be found in Appendix A. Proposition 1 tells that if we have an absolute perception coherence between the two models at a point $x$, the coherence level is equal to 1 at $x$. Now, a natural question to ask is whether the reverse statement holds. In short, our theoretical analysis yields a positive answer to this question. This will be presented in Section 3.3.
@@
Definition 3 (Local $\alpha$-perception coherence).
@@
Definition 4 (Global $\alpha$-perception coherence).

### v2

@@ Abstract
In this paper, we propose a method for transferring feature representation to lightweight student models from larger teacher models. We mathematically define a new notion called perception coherence. Based on this notion, we propose a loss function, which takes into account the dissimilarities between data points in feature space through their ranking. At a high level, by minimizing this loss function, the student model learns to mimic how the teacher model perceives inputs. More precisely, our method is motivated by the fact that the representational capacity of the student model is weaker than the teacher model. Hence, we aim to develop a new method allowing for a better relaxation. This means that, the student model does not need to preserve the absolute geometry of the teacher one, while preserving global coherence through dissimilarity ranking. Importantly, while rankings are defined only on finite sets, our notion of perception coherence extends them into a probabilistic form. This formulation depends on the input distribution and applies to general dissimilarity metrics. Our theoretical insights provide a probabilistic perspective on the process of feature representation transfer. Our experiments results show that our method outperforms or achieves on-par performance compared to strong baseline methods for representation transferring.
@@ 1 Introduction
The need of lightweight models and Knowledge distilling (KD). Deep learning models are at the forefront of performance in different domains and tasks, such as classification (He et al., 2016; Tan & Le, 2019; 2021; He et al., 2020) or object detection (Ren et al., 2015; Lin et al., 2017a; b; He et al., 2020). However, as the performance of neural networks increases, so does the cost of increasingly large models. In many applications where resources are limited (e.g. mobile devices) or one needs a fast execution, it is preferable to use lightweight models. Among the many approaches to obtain lightweight models, KD (Buciluǎ et al., 2006; Hinton et al., 2015; Wang & Yoon, 2021) stands out as a particularly interesting direction. A KD setup essentially consists of a teacher model and a student model. The student model (here the targeted lightweight model) is generally smaller than the teacher model, and therefore more efficient for the task in question. At a high level, the KD process consists in teaching the student model so that it mimics the way the teacher model understands the input. This is usually achieved by using a so-called transfer dataset. The student model is trained to behave like the teacher model on this set in some way . Many previous works have shown that KD improves the generalization capacity of the student model (Hinton et al., 2015; Chen et al., 2017; Zhu et al., 2018; Beyer et al., 2022; Huang et al., 2022). Different up-to-date KD approaches will be discussed in Section 2.
Figure 1: Illustration of feature representation transfer. A non-labeled transfer dataset is passed through both teacher and student models to obtain their respective feature representations. The transfer process consists in training the student model to somehow capture the input–input relationships encoded in the teacher model’s feature space.
@@
Feature representation transferring. Classical KD techniques generally focus on matching the outputs of student and teacher models, typically for classification tasks (Hinton et al., 2015). However, this approach has limitations, such as requiring the same number of classes between the teacher and student models. To address these issues, previous works (Passalis & Tefas, 2018a; Passalis et al., 2020b; Tian et al., 2020; Park et al., 2019) propose to directly transfer the feature space representations through learning. This enables the student model to better capture the input-input relations represented in the feature space of the teacher model, by preserving its geometry and distribution (illustrated in Fig. 1). This leads to more effective knowledge transfer. Furthermore, note that this approach is an unsupervised transferring method, as it only uses a transfer set without labels. As a consequence, the learned features can be used for various downstream tasks. Because of these appealing properties, our method aligns with this last approach.
@@
Motivation and an overview of our method. Generally, the student model is much smaller than the teacher model. Thus, in terms of representational power, the student model cannot learn to produce features with the exact same geometry as the teacher model. Hence, we aim to develop a method such that (1) the student model does not need to copy the feature representation of the teacher model; (2) the learned representation retains a certain overall consistency with the teacher model’s feature representation and (3) the method works for representation transferring between spaces having different dimensions. For these purposes, we develop a novel fairly simple notion called perception coherence, that is explained subsequently. To begin with, let us consider a reference input $x$ and a set of $n$ point $\{x_{i}\}_{i}^{n}$ (called compared set). These points are fed into the student and teacher models and are mapped into the feature space of each model. The feature space reflects how the models perceive the inputs. Placing ourselves in the two feature spaces, let $d_{1}$ and $d_{2}$ be dissimilarity measures for the teacher and student models, respectively. Then, if $d_{1}(x,x_{i})\leq d_{1}(x,x_{j})$ for some $i,j\in\{1,2,\dotsc,n\}$ ($i\neq j$), we expect that $d_{2}(x,x_{i})\leq d_{2}(x,x_{j})$ also holds. Less formally, if the teacher model perceives $x$ to be more similar to $x_{i}$ than to $x_{j}$, we expect the student to have the same perception. We refer to this as the property of perception coherence. By adopting this behavior for different reference points and compared sets, the student model learns to have a coherent perception with respect to the teacher model. Moreover, since such rankings are only defined over finite sets, we further generalize this concept into a probabilistic form—our formal definition of perception coherence (Definition 3.2)—which depends on the input distribution and applies to general dissimilarity metrics.
@@
Figure 2: General scheme of our method. In each mini-batch, we use each input as the reference point to compute the dissimilarity to the others (compared set). Then, using loss function in Eq. (4), the student model is trained to respect the perception coherence. Notice that $f_{1}$ and $f_{2}$ can be the whole model (when using the output layer) or a part of the model (when using an intermediate layer).
The general idea of our method is illustrated in Figure 2 (further details will be illustrated in Figure 3). Note that we do not quantify how much more similar $x_{i}$ is to $x$ than $x_{j}$. Instead, we only preserve the relative ranking of dissimilarities. This relaxation provides flexibility, as the student model is not required to replicate the exact geometry or distribution of the teacher model’s feature space. However, the ranking operation is non-differentiable. To address this, we introduce a simple differentiable approximation based on a soft ranking function (Eq. (3)). Our contribution can be summarized as follows:
  • We propose a new method for distilling knowledge to lightweight models, based on deep representation transferring. Central to our approach is a novel probabilistic concept, perception coherence, which naturally leads to a simple and easy-to-implement loss function (Eq. (4)) based on a soft ranking operation.
  • We provide theoretical insights through a probabilistic framework that connects representation transfer to local and global dissimilarity rankings in the feature spaces of teacher and student models. In particular, we show in Section 4.1 that the perception coherence level can be approximated via ranking operations on mini-batches with convergence rate $O(1/\sqrt{B})$, where $B$ is the mini-batch size. Furthermore, we shall also demonstrate that a larger global expected coherence level leads to a higher chance that the learned representation respects the dissimilarity ranking of the teacher one (Sections 4.2 and 4.3).
  • We conduct experiments on retrieval and classification tasks to compare our method with strong baseline methods. The results show that our method outperforms or achieves competitive performance with strong baseline methods on lightweight representation transferring settings. Our code is available at anonymized-code.
@@ 2 Related Work
KD based on soft label produced by teacher model. This approach has been studied since the early days of neural networks (Buciluǎ et al., 2006). Therein, the authors show that knowledge can be transferred from a large ensemble of models to a single small model. A standard KD method consists in using output of classification models (Hinton et al., 2015; Tang et al., 2016; Huang et al., 2022), with a temperature, leading to smoother soft labels. Based on the same spirit, (Chan et al., 2015) proposes to match the output distributions using Kullback-Leibler divergence. Soft label was also used in the context of domain adaption (Tzeng et al., 2015). While all these methods show the benefits of soft labels, the main limit of this approach is the need of having the same number of classes between models.
@@
KD based on distance matching. In Romero et al. (2014), the authors propose to minimize the distance between student and teacher’s features to guide the student model during training. However, in the case of dimension mismatch between the models, one need to pass through a linear transformation, which leads to substantial information loss. In Yim et al. (2017), the authors propose to minimize distance of the so-call FSP matrix between teacher and student models (capturing inter-layer relations). More advanced methods consist in minimizing the distance between the derivatives w.r.t. to the features (Czarnecki et al., 2017) or between the attention map of student and teacher (Zagoruyko & Komodakis, 2016). All these approaches require the dimension match, and need some form of interpolation in the case of dimension mismatch. Overcoming the drawback of dimension mismatch, Passalis & Tefas (2018b) proposes to match the the pair-wise similarity matrix of the training points (in the transfer set) between the student and teacher models. The method proposed in Yu et al. (2019) is to match the distance between pairs of samples. However, this brute force approach is not likely to work as the student model is much smaller than the teacher one (as discussed in Passalis et al. (2020b)).
@@
Feature representation transfer. Some previous works treat the representation transfer as the problem of distribution matching, based on Maximum Mean Discrepancy (Huang & Wang, 2017) or adversarial training (Belagiannis et al., 2018). The latter method is inspired by Generative Adversarial Networks (GANs) Goodfellow et al. (2014). One major drawback of distribution matching is the need of having the same dimension between the two feature spaces. Probabilistic frameworks based on mutual information are proposed in Ahn et al. (2019); Tian et al. (2020). However, the method of Ahn et al. (2019) requires a Gaussian assumption (this is not necessarily satisfied in practice) while the method of Tian et al. (2020) requires the input features to have the same dimension (hence requiring a linear transformation, which may reduce the useful information). Overcoming the requirement of dimension match, Park et al. (2019) proposes to transfer the input relations using distance-wise and angle-wise distillation loss. However, the structural information is quite limited, as it only considers a single triple of points at a time. Overcoming this drawback, Passalis & Tefas (2018a) proposes a method of probabilistic knowledge transfer (PKT). The high level idea of this last method is to transfer the local geometry learned by the teacher model based on kernel techniques. Based on this work, Passalis et al. (2020a; b) basically share the same idea, but they apply the technique for multiples intermediate layers with an auxiliary model. The benefits of using auxiliary models are discussed in Mirzadeh et al. (2020). As in Passalis & Tefas (2018a); Passalis et al. (2020a; b), our method takes into account the relation between the inputs and allows representation transfer between space having different dimensions. However, in Passalis & Tefas (2018a); Passalis et al. (2020a; b), a common kernel is used for all points (e.g. a Gaussian one with a fixed bandwidth), which cannot adapt to local density variation. This contrasts with our method, where we only rank the dissimilarities but do not use the dissimilarity magnitude. Thus, by using the dissimilarity rankings (or equivalently cumulative function of dissimilarities), our method captures local probabilistic closeness — a property that raw distance magnitudes fail to reflect in the presence of density variations.
@@ 3.1 Perception coherence and other notions
At a high level, if $f_{1}$ perceives $x$ to be more similar to $x_{1}$ than to $x_{2}$, then it should be also the case for $f_{2}$ after the representation transferring. This is represented through the feature space of each model, endowed with their corresponding dissimilarity metric. We present this intuition via a more formal concept named absolute perception coherence as follows.
Definition 3.1 (Absolute perception coherence)
@@
$F_{1}(x,x^{\prime}):=\mathbb{P}_{X}\left(d_{1}(x,X)\leq d_{1}(x,x^{\prime})\right),\;F_{2}(x,x^{\prime}):=\mathbb{P}_{X}(d_{2}(x,X)\leq d_{2}(x,x^{\prime}))\ .$ (1)
Remark 3.1
@@
Remark 3.2 (Relation between cumulative function and ranking)
Let us consider a discrete distribution uniformly distributed over $N$ examples $\{x_{i}\}_{i=1}^{N}$ and the extractor $f$ endowed with a dissimilarity metric $d$. We consider $x_{i}$ as the reference point and we compute dissimilarity between $x_{i}$ and other points $x_{j}$’s to obtain $D:=\{d(x_{i},x_{j})\}_{j=1}^{N}$. Assuming there is no ties, i.e., $d(x_{i},x_{k})\neq d(x_{i},x_{j})$ for all $k\neq j$, then, the rank of $d(x_{i},x_{j})$ in the set $D$ can be computed as $r(d_{ij}):=\sum_{k}\mathds{1}_{\{d_{ik}\leq d_{ij}\}}$. Moreover, the cumulative function $F(x_{i},x_{j})$ is the normalized rank of $d(x_{i},x_{j})$ in the set $D$. That is, $F(x_{i},x_{j})=r(d_{ij})/N=\frac{1}{N}\sum_{k}\mathds{1}_{\{d_{ik}\leq d_{ij}\}}$.
This remark allows us to intuitively understand the dissimilarity cumulative function through the lens of ranking. Hence, for a given reference point, we can know its “probabilistic distance” to the other points by using dissimilarity rankings to compute the cumulative function. This is illustrated in Fig. 3.
Figure 3: Illustration of dissimilarity ranking (in feature space) in the case of 4 points. Each of all the data points is considered as the reference point, and the dissimilarities to other data points are ranked to capture the cumulative distribution of dissimilarity (in feature space).
Using the cumulative function of Eq. (1), we define the perception coherence level as follows.
Definition 3.2 (Perception coherence level)
@@
$\phi_{f_{1},f_{2}}(x):=1-\mathbb{E}_{X}\left[\left|F_{1}(x,X)-F_{2}(x,X)\right|\right].$ (2)
Remark 3.3
@@
Proof 3.1
By definition, we have that $F_{1}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)$ and $F_{2}(x,X)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$. As $f_{2}$ is absolutely perception coherent with $f_{1}$ at $x$, we obtain that $d_{1}(x,X^{\prime})\leq d_{1}(x,X)\Leftrightarrow d_{2}(x,X^{\prime})\leq d_{2}(x,X)$. Thus, $\mathbb{P}_{X^{\prime}}\left(d_{1}(x,X^{\prime})\leq d_{1}(x,X)\right)=\mathbb{P}_{X^{\prime}}\left(d_{2}(x,X^{\prime})\leq d_{2}(x,X)\right)$, which leads to the result.
Proposition 3.1 tells that if we have an absolute perception coherence between the two models at a point $x$, the coherence level is equal to 1 at $x$. Now, a natural question to ask is whether the reverse statement holds. In short, our theoretical analysis yields a positive answer to this question. This will be presented in Section 4.
@@
Definition 3.3 (Local $\alpha$-perception coherence)
@@
Definition 3.4 (Global $\alpha$-perception coherence)